# TD3Net: A Temporal Densely Connected Multi-Dilated Convolutional Network for Lipreading

### v1

@@ Abstract
The word-level lipreading approach typically employs a two-stage framework with separate frontend and backend architectures to model dynamic lip movements. Each component has been extensively studied, and in the backend architecture, temporal convolutional networks (TCNs) have been widely adopted in state-of-the-art methods. Recently, dense skip connections have been introduced in TCNs to mitigate the limited density of the receptive field, thereby improving the modeling of complex temporal representations. However, their performance remains constrained owing to potential information loss regarding the continuous nature of lip movements, caused by blind spots in the receptive field. To address this limitation, we propose TD3Net, a temporal densely connected multi-dilated convolutional network that combines dense skip connections and multi-dilated temporal convolutions as the backend architecture. TD3Net covers a wide and dense receptive field without blind spots by applying different dilation factors to skip-connected features. Experimental results on a word-level lipreading task using two large publicly available datasets, Lip Reading in the Wild (LRW) and LRW-1000, indicate that the proposed method achieves performance comparable to state-of-the-art methods. It achieved higher accuracy with fewer parameters and lower floating-point operations compared to existing TCN-based backend architectures. Moreover, visualization results suggest that our approach effectively utilizes diverse temporal features while preserving temporal continuity, presenting notable advantages in lipreading systems. The code is available at our GitHub repository.
@@
Lipreading , Temporal Convolution , Dense Connectivity , Multi-dilation , TD3Net
@@ 3.2 Proposed Method
where $F_{l}(\cdot)$ indicates the following $l$th sequential operations: batch normalization (BN), followed by a 2D convolutional layer and nonlinear activation functions. $[x_{0};x_{1};\cdots;x_{l-1}]$ refers to the channel-wise concatenation of the input($x_{0}$) and feature maps received from $11,\cdots,l-$ 1 layers. Motivated by the aforementioned structures, Ma et al. [15] proposed the DC-TCN, which combines dense skip connections with TCN. In the TC layers of a dense block, the dilation factors $d$ are gradually increased as the layers are stacked (e.g., $d$ = 1, 2, 4, 8). However, the integration of the TC layers into DenseNet in a naive manner gives rise to blind spots in the receptive field [23].
@@
where $X_{l}=[x_{0};x_{1};\cdots;x_{l-1}]$ is the input for $l>0$, and ${d}_{i}\mathrm{TC}^{i}_{l}(\cdot)$ refers to the subset of the $H_{l}(\cdot)$ applied to $x_{i}$ with a dilation factor $d_{i}=2^{i}$, producing feature maps with a shape $k\times T$, where $k$ and $T$ denote the growth rate and sequence length, respectively. As depicted in Fig. 3(b), each skip-connected feature $x_{i}$ is processed individually by a dedicated sub-TC layer of $H_{l}(\cdot)$, with the dilation factor determined by the depth $i$ of the input feature. The outputs from these sub-TC layers are aggregated through element-wise summation to form the output feature map, where the activation highlighted in green is computed based on the sampled values (marked by black dots).
@@
This multi-dilated design adjusts the dilation factor for each input feature map, ensuring that the sampling interval does not exceed the receptive field of the individual activations within that map. As a result, the receptive fields of the sampled activations are temporally aligned without gaps, allowing all activations in the output feature map to reference continuous input regions without introducing blind spots. Consequently, the multi-dilated TC layers effectively capture temporal patterns at multiple scales in parallel—smaller dilations capture local dependencies, and larger ones capture long-range dependencies—while maintaining a seamless temporal flow of information.
To model temporal dependencies more comprehensively and reuse features across blocks efficiently, multiple TD2 blocks are densely connected, as illustrated in Fig. 4. We define this architecture as the TD3 block, and the TD3Net architecture is composed based on these TD3 blocks. Furthermore, to mitigate the indiscriminate increase in the number of feature maps resulting from the nested dense skip connections in the TD3 block, we apply a bottleneck layer comprising a 1D convolution with a kernel size of 1.
This layer is designed to transform the input and output of each TD2 block and the output from each TD3 block. The output channels of the TD3 and TD2 blocks are reduced by a compression ratio $c$ and transition ratio $t$, respectively $(0<c,t<1)$. In our experiments, the first layer, serving as a bottleneck for each TD2 block, was set to produce $4*k$ feature maps. Notably, TD3Net has several essential parameters, including the number of TD3 blocks $B$, number of TD2 blocks $N$, number of multi-dilated TC layers in a TD2 block $L$, growth rate $k$, and output channel reduction ratios $c$ and $t$. The steps for determining the optimal parameters of the network are described in detail in Sections 3.3 and 5.1. Furthermore, we provide PyTorch-style pseudocode in A to clarify the implementation of TD3Net.
@@ 3.3 Architecture Details
Our network begins with a $3{\times}7{\times}7$ convolutional layer (frame×$\times$×height×$\times$×width), followed by batch normalization (BN), ReLU activation, and a $1{\times}3{\times}3$ max pooling layer. This sequence replaces the first convolution and max pooling operations in the standard ResNet-18. Then, global average pooling is applied at the end of the last convolution, resulting in feature maps with a shape of $512\times T$ for inputs of shape $1\times T\times H\times W$, where the leading one indicates a single-channel grayscale image and $T$ (temporal length) is preserved throughout training. The backend architecture comprises sequential TD3 blocks, each containing an equal number of TD2 blocks. Each TD2 block includes the same number of multi-dilated TC layers with a filter size of 3. Each multi-dilated TC layer consists of sub-TC layers, where each sub-TC layer comprises a series of TC layers with batch normalization and ReLU activation. Similar to the transition layer in DenseNet-BC [22], the reduction ratios for output channels $c$ and $t$ are set to 0.5. No channel reduction is applied to the outputs of the last TD3 block.
@@ 4.1.1 LRW
LRW is the largest publicly available English word-level dataset and has been widely used in research. The dataset comprises 500 words spoken by more than 1000 speakers, divided into sub-datasets (training, validation, and evaluation sets), comprising 488K, 25K, and 25K videos, respectively. The training set includes 800–1000 utterances for each word, while the validation and evaluation sets contain 50 utterances each. All videos are 1.16 s long (29 frames at a frame rate of 25 fps) and have a fixed resolution of 256$\times$256 pixels.
In our study, we followed the data preprocessing approach used in previous studies [15, 14], and Fig. 5 illustrates an example of the process. First, we employed the pretrained facial landmark detector from the Dlib toolkit [49] to locate 68 key points on the face, including those on the eyes, eyebrows, nose, and mouth. The extracted landmark coordinates enabled us to center the face in each image by aligning it based on these coordinates. Centering the face enhances the performance of lipreading models by emphasizing lip movements, maintaining consistency across frames, and reducing noise by excluding extraneous visual information. Subsequently, we cropped the mouth region from the aligned image to a size of 96$\times$96 pixels and converted it into a grayscale image. This conversion enhanced the robustness of the model to variations in lighting and color while also reducing computational complexity.
@@ 4.1.2 LRW-1000
The data were preprocessed according to the pipelines introduced by Feng et al. [9]. To ensure consistent video lengths, as in the LRW dataset, all videos were adjusted to 40 frames using zero padding. Then, each frame, which included an image of the mouth region, was resized to 96$\times$96 pixels and transformed into a grayscale image.
Table 1: Performance comparison with state-of-the-art word-level lipreading methods on the LRW and LRW-1000 datasets. $WB$ denotes methods using word boundaries as binary indicator vectors. Means and standard deviations are reported over three runs.
Method Backbone Accuracy (%) $WB$
@@
Zhao et al. [24] 3D $+$ ResNet BiGRU 84.41 38.79
Zhang et al. [10] / Cutout 3D $+$ ResNet BiGRU 85.02 45.24
Martinez et al. [14] 3D $+$ ResNet MS-TCN 85.30 41.4
Feng et al. [9] / Training tricks 3D $+$ SE-ResNet BiGRU 88.40 55.70
Ma et al. [4] / KD & Ensemble 3D $+$ ResNet MS-TCN 88.50$\dagger$ 46.60$\dagger$
Kim et al. [6] / Cross-modal memory 3D $+$ ResNet BiGRU 85.40* 50.82*
Ma et al. [15] 3D $+$ ResNet DC-TCN 88.36 43.65
Kim et al. [7] / Multi-head memory 3D $+$ ResNet MS-TCN 88.50* 53.82*
@@
89.57$\dagger$ -
TD3Net-Best (ours) 3D + EfficientNetV2-S TD3Net 89.86±$\pm 0.06$ 52.01±$\pm 0.10$
3D + ResNet 89.54±$\pm 0.08$ 51.41±$\pm 0.07$
3D + ResNet 91.41±$\pm 0.10$ 59.20±$\pm 0.11$
@@ 4.2 Training Details
We followed a training procedure similar to that described in previous studies [15]. The training settings were nearly identical for both the LRW and LRW-1000 datasets, with specific adjustments made to prevent overfitting. A network with randomly initialized weights was trained to recognize words in the video by minimizing the cross-entropy error. For training, we set a batch size of 32 and selected 120 epochs. We used the AdamW optimizer [54] with a $l_{2}$ weight decay of 1 ×$\times 10^{-2}$. The learning rate was initialized at 3 ×$\times 10^{-4}$ and steadily decreased to 5 ×$\times 10^{-8}$ using a cosine annealing scheduler. The images of the mouth region were randomly cropped to a size of $88{\times}8$ 88, and a horizontal flip was applied with a probability of 0.5 for data augmentation. Referring to the code released by Ma et al. [15], we also included the Mixup [55] data augmentation technique by setting the value of $\alpha$ to 0.4, where $\alpha$ determined the degree of interpolation between video-word pairs. A dropout [56] with a probability of 0.2 for LRW and 0.1 for LRW-1000 was applied to avoid overfitting after each TC layer, which comprised a 1D dilated convolutional layer, BN, and ReLU activation function. For testing, each frame of the video was cropped from the center to $88{\times}8$ 88 pixels. All experimental results are reported as the mean and standard deviation measured over three runs conducted in the following environments: UBUNTU 20.04, PYTHON 3.8.13, PyTorch 1.8.0, CUDA 11.1, and NVIDIA RTX 3090.
@@ 5 Discussion
4 linear 25.87 88.28±$\pm 0.14$
exponential 88.56±$\pm 0.09$
5 linear 30.75 (Base) 89.03±$\pm 0.06$
exponential 89.36±$\pm 0.09$
6 linear 36.27 89.15±$\pm 0.13$
exponential 89.37±$\pm 0.09$
@@ 5.2 Ablation Study
This section compares three configurations to analyze how dilation strategies applied to skip connections influence temporal modeling, through both quantitative and qualitative evaluations. Specifically, we investigated the impact of blind spots that may arise from integrating dilation into skip connections and assessed whether the proposed multi-dilated TC layers can effectively mitigate these issues. To this end, the multi-dilated TC layers were replaced with non-dilated TC layers and with standard TC layers that use a single dilation factor, which may introduce blind spots. All configurations were designed with comparable parameter counts to ensure a fair comparison. Experimental results are summarized in Fig. 6 and Table 6, with configurations (B) and (C) corresponding to graphs (a) and (b) in Fig. 3, respectively.
@@
(A) w/o dilation 30.75 87.83±$\pm 0.10$
(B) Standard dilation 30.75 88.62±$\pm 0.13$
(C) Multi dilation (ours) 30.75 89.36±$\pm 0.09$

### v4

@@ Abstract
The word-level lipreading approach typically employs a two-stage framework with separate frontend and backend architectures to model dynamic lip movements. Each component has been extensively studied, and in the backend architecture, temporal convolutional networks (TCNs) have been widely adopted in state-of-the-art methods. Recently, dense skip connections have been introduced in TCNs to mitigate the limited density of the receptive field, thereby improving the modeling of complex temporal representations. However, their performance remains constrained owing to potential information loss regarding the continuous nature of lip movements, caused by blind spots in the receptive field. To address this limitation, we propose TD3Net, a temporal densely connected multi-dilated convolutional network that combines dense skip connections and multi-dilated temporal convolutions as the backend architecture. TD3Net covers a wide and dense receptive field without blind spots by applying different dilation factors to skip-connected features. Experimental results on a word-level lipreading task using two large publicly available datasets, Lip Reading in the Wild (LRW) and LRW-1000, indicate that the proposed method achieves performance comparable to state-of-the-art methods. It achieved higher accuracy with fewer parameters and lower floating-point operations compared to existing TCN-based backend architectures. Moreover, visualization results suggest that our approach effectively utilizes diverse temporal features while preserving temporal continuity, presenting notable advantages in lipreading systems. The code is available at our GitHub repository (https://github.com/Leebh-kor/TD3Net).
@@
Visual speech recognition , Lipreading , Temporal convolution , Dense connectivity , Multi-dilation
@@ 2.2.2 Backend – Temporal Model
2.2.3 Audio-Visual Speech Recognition
Audio-visual speech recognition (AVSR) has emerged as a prominent research direction, aiming to leverage both visual and audio modalities for more robust speech understanding. This line of research builds upon the progress made in both automatic speech recognition (ASR), which primarily focuses on audio signals, and visual-only lipreading, which seeks to decode speech from lip movements. Building on this foundation, numerous AVSR approaches [40, 43, 44, 45, 46, 47] have been proposed in recent years, leveraging advances in visual-auditory representation learning.
From an architectural perspective, most AVSR systems still adopt visual processing pipelines similar to those used in conventional lipreading. Specifically, visual features are typically extracted through a frontend encoder, often based on ResNet variants, which are then passed to a temporal backend model that captures sequential dependencies across video frames. Among various backend architectures, transformer-based models are frequently employed owing to their ability to model long-range temporal dependencies effectively. Additionally, Conformer [42] and Branchformer [48] variants have been introduced to improve the modeling of both local and global temporal features, further enhancing the generalization capabilities of audio-visual systems.
Although our study focuses solely on visual input, the proposed TD3Net, an enhanced architecture based on TCN, has the potential to serve as an efficient backend module in multimodal lipreading systems. While transformer-based temporal encoders are widely used for their scalability and ability to model long-range dependencies, their self-attention mechanism incurs quadratic computational complexity with respect to sequence length. This limitation may hinder their practicality in real-time or low-resource scenarios, particularly when processing long video sequences. By contrast, the proposed TD3Net is inherently based on the TCN architecture, characterized by low computational complexity and a causal design, enabling both low-latency inference and real-time applicability. These properties make it a strong candidate not only as a lightweight alternative but also as a complementary module to transformer-based backend architectures. Accordingly, the architectural design of TD3Net indicates its broader applicability to future multimodal lipreading frameworks, where efficient temporal modeling remains a critical component.
@@ 3.2 Proposed Method
where $F_{l}(\cdot)$ indicates the following $l$th sequential operations: batch normalization (BN), followed by a 2D convolutional layer and nonlinear activation functions. $x_{0};x_{1};\cdots;x_{l-1}]$ refers to the channel-wise concatenation of the input($x_{0}$) and feature maps received from $1,\cdots,l-1$ layers. Motivated by the aforementioned structures, Ma et al. [15] proposed the DC-TCN, which combines dense skip connections with TCN. In the TC layers of a dense block, the dilation factors $d$ are gradually increased as the layers are stacked (e.g., $d$ = 1, 2, 4, 8). However, the integration of the TC layers into DenseNet in a naive manner gives rise to blind spots in the receptive field [23].
@@
where $X_{l}=[x_{0};x_{1};\cdots;x_{l-1}]$ is the input for $l>0$, and ${d}_{i}\mathrm{TC}^{i}_{l}(\cdot)$ refers to the subset of the $H_{l}(\cdot)$ applied to $x_{i}$ with a dilation factor $d_{i}=2^{i}$, producing feature maps with a shape $k\times T$, where $k$ and $T$ denote the growth rate and sequence length, respectively.
As depicted in Fig. 3(b), each skip-connected feature $x_{i}$ is processed individually by a dedicated sub-TC layer of $H_{l}(\cdot)$, with the dilation factor determined by the depth $i$ of the input feature. The outputs from these sub-TC layers are aggregated through element-wise summation to form the output feature map, where the activation highlighted in green is computed based on the sampled values (marked by black dots).
@@
This multi-dilated design adjusts the dilation factor for each input feature map, ensuring that the sampling interval does not exceed the receptive field of the individual activations within that map. As a result, the receptive fields of the sampled activations are temporally aligned without gaps, allowing all activations in the output feature map to reference continuous input regions without introducing blind spots. Consequently, the multi-dilated TC layers effectively capture temporal patterns at multiple scales in parallel—smaller dilations capture local dependencies, and larger ones capture long-range dependencies—while maintaining a seamless temporal flow of information. To model temporal dependencies more comprehensively and reuse features across blocks efficiently, multiple TD2 blocks are densely connected, as illustrated in Fig. 4. We define this architecture as the TD3 block, and the TD3Net architecture is composed based on these TD3 blocks.
Furthermore, to mitigate the indiscriminate increase in the number of feature maps resulting from the nested dense skip connections in the TD3 block, we apply a bottleneck layer comprising a 1D convolution with a kernel size of 1. This layer is designed to transform the input and output of each TD2 block and the output from each TD3 block. The output channels of the TD3 and TD2 blocks are reduced by a compression ratio $c$ and transition ratio $t$, respectively $(0<c,t<1)$. In our experiments, the first layer, serving as a bottleneck for each TD2 block, was set to produce 4k4*k feature maps. Notably, TD3Net has several essential parameters, including the number of TD3 blocks $B$, number of TD2 blocks $N$, number of multi-dilated TC layers in a TD2 block $L$, growth rate $k$, and output channel reduction ratios $c$ and $t$. The steps for determining the optimal parameters of the network are described in detail in Sections 3.3 and 5.1. Furthermore, we provide PyTorch-style pseudocode in A to clarify the implementation of TD3Net.
@@ 3.3 Architecture Details
[…]
@@ 4.1.1 LRW
LRW is the largest publicly available English word-level dataset and has been widely used in research. The dataset comprises 500 words spoken by more than 1000 speakers, divided into sub-datasets (training, validation, and evaluation sets), comprising 488K, 25K, and 25K videos, respectively. The training set includes 800–1000 utterances for each word, while the validation and evaluation sets contain 50 utterances each. All videos are 1.16 s long (29 frames at a frame rate of 25 fps) and have a fixed resolution of $\times256$ pixels.
In our study, we followed the data preprocessing approach used in previous studies [15, 14], and Fig. 5 illustrates an example of the process. First, we employed the pretrained facial landmark detector from the Dlib toolkit [55] to locate 68 key points on the face, including those on the eyes, eyebrows, nose, and mouth. The extracted landmark coordinates enabled us to center the face in each image by aligning it based on these coordinates. Centering the face enhances the performance of lipreading models by emphasizing lip movements, maintaining consistency across frames, and reducing noise by excluding extraneous visual information. Subsequently, we cropped the mouth region from the aligned image to a size of $\times96$ pixels and converted it into a grayscale image. This conversion enhanced the robustness of the model to variations in lighting and color while also reducing computational complexity.
@@ 4.1.2 LRW-1000
The data were preprocessed according to the pipelines introduced by Feng et al. [9]. To ensure consistent video lengths, as in the LRW dataset, all videos were adjusted to 40 frames using zero padding. Then, each frame, which included an image of the mouth region, was resized to $\times96$ pixels and transformed into a grayscale image.
Table 1: Performance comparison with state-of-the-art word-level lipreading methods on the LRW and LRW-1000 datasets. WBWB denotes methods using word boundaries as binary indicator vectors. Means and standard deviations are reported over three runs.
Method Backbone Accuracy (%) WBWB
@@
Zhao et al. [24] 3D ++ ResNet BiGRU 84.41 38.79
Zhang et al. [10] / Cutout 3D ++ ResNet BiGRU 85.02 45.24
Martinez et al. [14] 3D ++ ResNet MS-TCN 85.30 41.4
Feng et al. [9] / Training tricks 3D ++ SE-ResNet BiGRU 88.40 55.70
[…]
Kim et al. [6] / Cross-modal memory 3D ++ ResNet BiGRU 85.40* 50.82*
Ma et al. [15] 3D ++ ResNet DC-TCN 88.36 43.65
Kim et al. [7] / Multi-head memory 3D ++ ResNet MS-TCN 88.50* 53.82*
@@
[…]
@@ 4.2 Training Details
We followed a training procedure similar to that described in previous studies [15]. The training settings were nearly identical for both the LRW and LRW-1000 datasets, with specific adjustments made to prevent overfitting. A network with randomly initialized weights was trained to recognize words in the video by minimizing the cross-entropy error. For training, we set a batch size of 32 and selected 120 epochs. We used the AdamW optimizer [60] with a $l_{2}$ weight decay of 1 $\times 10^{-2}$. The learning rate was initialized at 3 $\times 10^{-4}$ and steadily decreased to 5 $\times 10^{-8}$ using a cosine annealing scheduler. The images of the mouth region were randomly cropped to a size of $88{\times}88$, and a horizontal flip was applied with a probability of 0.5 for data augmentation. Referring to the code released by Ma et al. [15], we also included the Mixup [61] data augmentation technique by setting the value of $\alpha$ to 0.4, where $\alpha$ determined the degree of interpolation between video-word pairs. A dropout [62] with a probability of 0.2 for LRW and 0.1 for LRW-1000 was applied to avoid overfitting after each TC layer, which comprised a 1D dilated convolutional layer, BN, and ReLU activation function. For testing, each frame of the video was cropped from the center to $88{\times}88$ pixels. All experimental results are reported as the mean and standard deviation measured over three runs conducted in the following environments: UBUNTU 20.04, PYTHON 3.8.13, PyTorch 1.8.0, CUDA 11.1, and NVIDIA RTX 3090.
@@ 5 Discussion
[…]
@@ 5.2 Ablation Study
5.2.1 Limitations of TD2 Block in Isolation
To independently evaluate the effectiveness of the TD2 block, we conducted a comparative experiment by replacing the backend architecture of TD3Net with two alternative configurations:
  • Dense-TCN: A baseline model composed of four dense blocks, each densely connected with dilated temporal convolution (TC) layers. The block depths are 6, 12, 24, and 16 TC layers, respectively. Within each block, the dilation factor for the $i$-th TC layer, where $i$ starts from 0, follows the increasing pattern $2\times(i\bmod 8)+1$. The dense blocks in this model correspond to graph (a) in Fig. 3.
  • TD2Net: A TD3Net variant composed of four TD2 blocks, where each block is densely connected with multi-dilated TC layers. The only difference from Dense-TCN is the use of multi-dilated TC layers, with each layer assigned a dilation factor following the increasing pattern $2\times(i\bmod 8)+1$. The TD2 blocks in this model correspond to graph (b) in Fig. 3.
Table 6: Ablation results on the TD2 block structure evaluated on the LRW test dataset. The dilation factor of (A) is identical to the maximum dilation value of the corresponding layer from (B).
Configuration # Params (M) Accuracy (%)
[…]
Table 6 shows that, despite incorporating multi-dilation to capture richer temporal information, TD2Net performed worse than Dense-TCN. In TD2Net, each block applies distinct dilation factors to the skip-connected features, aiming to extract multi-scale temporal patterns in parallel.
However, the structural complexity introduced by aggregating multi-dilated features across most layers can destabilize gradient flow during early training, impairing the model’s ability to capture salient temporal cues. In practice, removing the dense connections between TD2 blocks in TD3Net-base resulted in unstable training and convergence failure. Training stability was restored only after restructuring the network into four TD2 blocks with progressively increasing depths, inspired by DenseNet-121 [22]. This configuration facilitates gradual feature expansion and encourages efficient feature reuse while maintaining manageable parameter growth.
By contrast, Dense-TCN, with standard dilated TC layers, maintains structural simplicity and more coherent gradient flow, naturally offering easier optimization. These results suggest that module-level (TD2 block) design alone is insufficient to fully realize the expressive capacity of the network. To efficiently exploit structurally complex and expressive components such as TD2 blocks, incorporating additional dense connections across blocks is crucial. Such higher-level connections promote stable temporal aggregation and robust feature reuse, essential for effective temporal representation learning.
5.2.2 Analysis of Multi-Dilation and TD2 Block Integration
This section compares three configurations to analyze how dilation strategies applied to skip connections influence temporal modeling, through both quantitative and qualitative evaluations. Specifically, we investigated the impact of blind spots that may arise from integrating dilation into skip connections and assessed whether the proposed multi-dilated TC layers can effectively mitigate these issues. To this end, the multi-dilated TC layers were replaced with non-dilated TC layers and with standard TC layers that use a single dilation factor, which may introduce blind spots. All configurations were designed with comparable parameter counts to ensure a fair comparison. Experimental results are summarized in Fig.6 and Table7, with configurations (B) and (C) corresponding to graphs (a) and (b) in Fig. 3, respectively. Furthermore, motivated by the structural complexity and training instability of TD2 blocks discussed in Section 5.2.1, we also investigated the necessity of dense connections between TD2 blocks. This analysis, informed by both previous and current experimental results, offers further insight into the nested dense connection design of TD3Net.
@@ 5.2.2 Analysis of Multi-Dilation and TD2 Block Integration
[…]
@@
Meanwhile, while configuration (B) is based on Dense-TCN’s dense block structure and configuration (C) on TD2Net’s TD2 block, the current results differ from previous experiments where Dense-TCN outperformed TD2Net. Configuration (C) outperformed (B), demonstrating that within a nested dense connection structure, multi-dilated layers are more effective than standard dilated layers. These findings highlight that effectively exploiting the expressive capacity of the TD2 block requires structural support via dense integration across blocks, thereby emphasizing the necessity of the TD3 block.
5.2.3 Summary of Findings
The experimental results confirm that multi-dilated layers within the TD2 block effectively mitigate blind spots induced by standard dilation, enabling more comprehensive temporal information capture. Despite the high representational capability of the TD2 block, it exhibits inherent limitations in terms of training stability and overall performance when used in isolation. These findings suggest that simply employing multi-dilated layers is not sufficient; careful architectural design—such as that of TD3Net—is essential to fully harness their potential. To address these limitations, TD3Net incorporates additional dense connections between TD2 blocks, enhancing temporal information integration. This nested integration approach substantially leverages the representational power of TD2 blocks and has been empirically shown to improve lipreading performance.
5.3 TD3Net for Real-Time Lipreading and AVSR Systems
Most existing lipreading studies have primarily focused on improving recognition accuracy, often overlooking the practical constraints associated with real-time deployment [63]. For example, recent studies tend to adopt computationally intensive temporal architectures such as transformer, Conformer [42], and Branchformer [48], which prioritize performance over inference efficiency. However, as lipreading systems advance toward real-world applications, the demand for real-time operation is expected to grow. In this regard, the proposed TD3Net is well-suited for such scenarios, as it is based on TCNs that offer several efficiency advantages, such as relatively low computational complexity and a causal design that is particularly beneficial for streaming or low-latency applications.
To illustrate the computational differences between temporal modeling architectures, consider a 1D convolution layer with a kernel size of $k=3$, which has a computational cost of $3NC^{2}$, where $N$ is the sequence length and $C$ is the channel dimension. By contrast, a self-attention layer includes three linear projections to compute the query, key, and value matrices, each with a cost of $NC^{2}$, and an attention operation with a cost of $N^{2}C$. The total computational cost of the self-attention layer is therefore approximately $3NC^{2}+N^{2}C$. This gap becomes especially problematic in sentence-level lipreading, where the input sequence is typically much longer than in word-level tasks. As $N$ increases, the quadratic term $N^{2}C$ dominates, leading to substantial overhead in real-time or resource-constrained environments. In such cases, lightweight alternatives such as TCN-based models offer a more practical and scalable solution for temporal modeling.
Furthermore, the use of standard self-attention in real-time lipreading scenarios presents an additional challenge related to causality. Most self-attention mechanisms are noncausal by default, meaning they attend to both past and future frames. While this is beneficial for offline recognition, it is incompatible with real-time inference, where future frames are unavailable. To enable streaming applications, causal (masked) self-attention must be employed. However, if the model is trained with full self-attention but deployed with causal masking at inference time, a significant training–inference mismatch occurs. This mismatch can lead to noticeable performance degradation, as the model learns to rely on future context during training that is no longer accessible at test time. By contrast, TD3Net is designed to support causal processing during both training and inference, enabling consistent behavior without training-inference mismatch. This property is essential for stable and reliable real-time lipreading.
This robustness in real-time settings makes TD3Net a promising candidate not only as a standalone temporal model but also as a complementary module in broader lipreading and audiovisual speech recognition (AVSR) systems. As mentioned above, recent lipreading and AVSR frameworks commonly employ Conformer and Branchformer architectures to jointly model global and local temporal dependencies through a combination of self-attention and convolution. However, their reliance on standard convolution blocks with fixed dilation restricts the diversity of temporal patterns that can be captured. Replacing these blocks with TD3Net’s multi-dilated convolution blocks could enhance temporal modeling by enabling the extraction of more diverse and flexible patterns across varying time scales, potentially improving performance in both visual-only and multimodal lipreading systems.